OpenStudy (anonymous):

An arc of length 30 meters is formed by a central angle A on a circle of radius 15. The measure of A in degrees (to two decimal places) is.... 114.59 - right?

OpenStudy (whpalmer4):

Radius of the circle is 15, so the circumference is \[C = 2\pi r = 2\pi(15) =30 \pi\]Arc length of \(30:30\pi\) subtended by angle \(A:360\) \[\frac{30}{30\pi} = \frac{A}{360}\]\[A = \frac{30*360}{30\pi} \approx 114.592\] so your answer is correct.

OpenStudy (anonymous):

Thank you!!